# Suppressing Beam Squint Effect For Near-Field Wideband Communication Through Movable Antennas

### v2

@@ I Introduction
II System Model and Problem Formulation
As shown in Fig. 1, we consider a wideband MISO communication system where a 2-D planar MA array with $M$ antennas residing in the $yoz$-plane serves a single-antenna user. Without loss of generality, we assume that the MA array has a square shape of size $A\times A$, locates its center at the origin of the coordinate system and arranges its neighboring edges parallel with the $y$- and $z$-axis, respectively. Denote $\mathcal{M}\triangleq\{1,\cdots,M\}$ as the set of MAs and $\bm{p}_{m}\triangleq[0,y_{m},z_{m}]^{\mathsf{T}},\;m\in\mathcal{M}$, as the position of the $m$-th MA. Since all MAs should maintain within the array surface area, the restriction $\bm{p}_{m}\in\mathcal{C}_{\mathrm{MA}}\triangleq\{(0,y,z)|y\in[-\frac{A}{2},\frac{A}{2}],\;z\in[-\frac{A}{2},\frac{A}{2}]\}$, for all $m\in\mathcal{M}$, should be satisfied with $\mathcal{C}_{\mathrm{MA}}$ indicating the feasible dwelling region of the MAs [4]-[7].
@@ II System Model and Problem Formulation
As indicated by (3), analog beamforming gain depends on the subcarrier frequency $f$. For the wideband system, the carrier frequency $f$ varies over a wide spectrum and hence incurs severe fluctuating beamforming gain across subcarriers, which is the so-called beam squint effect. Note that the squinted beam pattern is related to the path delay between each antenna and user [8]-[11], which can be affected by the positions of antennas. Hence, the beam squint effect could be alleviated via delicately configuring MAs’ positions, in other words, $\mathsf{g}(\{\bm{p}_{m}\}_{m=1}^{M},f_{\mathrm{c}},f)$ among different $f$’s can be improved. To provide more balanced communication rate among different subcarrier’s data stream [8], [9], we propose to maximize the minimum analog beamforming among the entire band. Specifically, denote the subcarrier frequencies as $\{f_{l}\}_{l=0}^{L}$ with $f_{0}<\cdots<f_{L}$. The configuration of MAs is formulated as the following problem
@@ III SV-Based Classical Solution
1:  Initialize feasible $\{\bm{t}_{m}^{(0)}\}_{m=1}^{M}$ and $n=0$;
2:  repeat
3:     for $m=1$ to $M$ do
4:        update $(\bm{t}_{m}^{(n+1)},\kappa^{(n+\frac{m}{M})})$ by solving $(\mathcal{P}4)$;
5:     end for
6:     $n:=n+1$;
7:  until convergence
@@ B. Update of $f$
Obviously, there exists a possibility that $f^{(n,i+1)}$ falls out of the feasible domain $[f_{0},f_{L}]$. When this occurs, we should project $f^{(n,i+1)}$ onto $[f_{0},f_{L}]$, that is:
@@ C. Update of $\bm{\alpha}_{m}$
1:  Initialize feasible $\{\bm{t}_{m}^{(0)}\}_{m=1}^{M}$ and $n=0$;
2:  repeat
3:     for $m=1$ to $M$ do
4:        set $\bm{t}_{m}^{(n,0)}:=\bm{t}_{m}^{(n)}$, $\bm{\alpha}_{m}^{(n,0)}:=\bm{t}_{m}^{(n,0)}$ and $i=0$;
5:        randomly generate $f^{(n,0)}\in[f_{0},f_{L}]$;
6:        repeat
7:           update $\bm{t}_{m}^{(n,i+1)}$ by (13);
8:           project $\bm{t}_{m}^{(n,i+1)}$ by solving $(\mathcal{P}6)$ (if needed);
9:           update $f^{(n,i+1)}$ by (16);
110:           project $f^{(n,i+1)}$ by $f^{(n,i+1)}=f_{0}$
or $f^{(n,i+1)}=f_{L}$ (if needed);
11:           update $\bm{\alpha}_{m}^{(n,i+1)}$ by (17);
12:           $i:=i+1$;
13:        until convergence
14:        set $\bm{t}_{m}^{(n+1)}:=\bm{t}_{m}^{(n,\infty)}$;
15:     end for
16:     $n:=n+1$;
17:  until convergence
@@ B. Derivation of a Lower Bound $\mathbf{M}_{m,l}$ for Hessian in (7)
×$\times\mathsf{max}\bigg{\{}\bigg{|}\bigg{[}\frac{\partial\mathsf{a}(\bm{t}_{m})}{\partial y_{m}}\frac{\partial\mathsf{a}(\bm{t}_{m})}{\partial z_{m}}\bigg{]}_{\!\mathrm{L}}\bigg{|},\bigg{|}\bigg{[}\frac{\partial\mathsf{a}(\bm{t}_{m})}{\partial y_{m}}\frac{\partial\mathsf{a}(\bm{t}_{m})}{\partial z_{m}}\bigg{]}_{\!\mathrm{U}}\bigg{|}\bigg{\}},$ (36)
where $[x]_{\mathrm{L}}$ and $[x]_{\mathrm{U}}$ denote the lower and upper bounds of scalar $x$, respectively. After some manipulations, both bounds in the second row of (36) can be written as

### v3

@@ I Introduction
II System Model and Problem Formulation
As shown in Fig. 1, we consider a wideband MISO communication system where a 2-D planar MA array with $M$ antennas residing in the yozyoz-plane serves a single-antenna user. Without loss of generality, we assume that the MA array has a square shape of size $A\times A$, locates its center at the origin of the coordinate system and arranges its neighboring edges parallel with the $y$- and $z$-axis, respectively. Denote $\mathcal{M}\triangleq\{1,\cdots,M\}$ as the set of MAs and $\bm{p}_{m}\triangleq[0,y_{m},z_{m}]^{\mathsf{T}},\;m\in\mathcal{M}$, as the position of the $m$-th MA. Since all MAs should maintain within the array surface area, the restriction $\bm{p}_{m}\in\mathcal{C}_{\mathrm{MA}}\triangleq\{(0,y,z)|y\in[-\frac{A}{2},\frac{A}{2}],\;z\in[-\frac{A}{2},\frac{A}{2}]\}$, for all $m\in\mathcal{M}$, should be satisfied with $\mathcal{C}_{\mathrm{MA}}$ indicating the feasible dwelling region of the MAs [4]-[7].
@@ II System Model and Problem Formulation
As indicated by (3), analog beamforming gain depends on the subcarrier frequency $f$. For the wideband system, the carrier frequency $f$ varies over a wide spectrum and hence incurs severe fluctuating beamforming gain across subcarriers, which is the so-called beam squint effect. Note that the squinted beam pattern is related to the path delay between each antenna and user [8]-[11], which can be affected by the positions of antennas. Hence, the beam squint effect could be alleviated via delicately configuring MAs’ positions, in other words, $\mathsf{g}(\{\bm{p}_{m}\}_{m=1}^{M},f_{\mathrm{c}},f)$ among different $f$’s can be improved. To provide more balanced communication rate among different subcarrier’s data stream [8], [9], [16], we propose to maximize the minimum analog beamforming among the entire band. Specifically, denote the subcarrier frequencies as $\{f_{l}\}_{l=0}^{L}$ with $f_{0}<\cdots<f_{L}$. The configuration of MAs is formulated as the following problem
@@ III SV-Based Classical Solution
1:  Initialize feasible $\{\bm{t}_{m}^{(0)}\}_{m=1}^{M}$ and $n=0$; 2:  repeat 3:   for $m=1$ to $M$ do 4:    update $(\bm{t}_{m}^{(n+1)},\kappa^{(n+\frac{m}{M})})$ by solving $(\mathcal{P}4)$; 5:   end for 6:   $n:=n+1$; 7:  until convergence
@@ B. Update of $f$
Obviously, there exists a possibility that $f^{(n,i+1)}$ falls out of the feasible domain $f_{0},f_{L}]$. When this occurs, we should project $f^{(n,i+1)}$ onto $f_{0},f_{L}]$, that is:
@@ C. Update of $\bm{\alpha}_{m}$
1:  Initialize feasible $\{\bm{t}_{m}^{(0)}\}_{m=1}^{M}$ and $n=0$; 2:  repeat 3:   for $m=1$ to $M$ do 4:    set $\bm{t}_{m}^{(n,0)}:=\bm{t}_{m}^{(n)}$, $\bm{\alpha}_{m}^{(n,0)}:=\bm{t}_{m}^{(n,0)}$ and $i=0$; 5:    randomly generate $f^{(n,0)}\in[f_{0},f_{L}]$; 6:    repeat 7:     update $\bm{t}_{m}^{(n,i+1)}$ by (13); 8:     project $\bm{t}_{m}^{(n,i+1)}$ by solving $(\mathcal{P}6)$ (if needed); 9:     update $f^{(n,i+1)}$ by (16); 110:     project $f^{(n,i+1)}$ by $f^{(n,i+1)}=f_{0}$ or $f^{(n,i+1)}=f_{L}$ (if needed); 11:     update $\bm{\alpha}_{m}^{(n,i+1)}$ by (17); 12:     $i:=i+1$; 13:    until convergence 14:    set $\bm{t}_{m}^{(n+1)}:=\bm{t}_{m}^{(n,\infty)}$; 15:   end for 16:   $n:=n+1$; 17:  until convergence
@@ VI Conclusion
Although this paper only considers a single-user scenario, the proposed scheme for tackling beam squint issue can be extended to multi-user systems, where other objectives, e.g., spectral efficiency, rather than array gain, may be considered due to multi-user interference. Besides, the latest proposed six-dimensional movable antenna (6DMA), which offers both movability and rotatability, could further enhance beam squint mitigation [20]. These open problems are worthy to be investigated in future researches.
@@ B. Derivation of a Lower Bound $\mathbf{M}_{m,l}$ for Hessian in (7)
$\times\mathsf{max}\bigg{\{}\bigg{|}\bigg{[}\frac{\partial\mathsf{a}(\bm{t}_{m})}{\partial y_{m}}\frac{\partial\mathsf{a}(\bm{t}_{m})}{\partial z_{m}}\bigg{]}_{\!\mathrm{L}}\bigg{|},\bigg{|}\bigg{[}\frac{\partial\mathsf{a}(\bm{t}_{m})}{\partial y_{m}}\frac{\partial\mathsf{a}(\bm{t}_{m})}{\partial z_{m}}\bigg{]}_{\!\mathrm{U}}\bigg{|}\bigg{\}},$ (36)
where $x]_{\mathrm{L}}$ and $x]_{\mathrm{U}}$ denote the lower and upper bounds of scalar $x$, respectively. After some manipulations, both bounds in the second row of (36) can be written as